I need help plz
A hydrocarbon X effuses through a pinhole 1.67 tim
Answer #1

Graham's Law :- The rates of effusion of two gases are inversely proportional to the square roots of their molar masses.

Therefore

\frac{Rate_1}{Rate_2}= \sqrt{\frac{M_2}{M_1}}

where:

Rate1 is the rate of effusion of the first gas (volume or number of moles per unit time).

Rate2 is the rate of effusion for the second gas.

M1 is the molar mass of gas 1

M2 is the molar mass of gas 2.

Or

M_1R_1^2=M_2R_2^2

Given is:-

Molor mass of Hydrocarbon X is = M_1

Molor mass of Kr is = 83.80 g/mol

Rate of effusion of Hydrocarbon X = 1.67

Rate of effusion of Kr = 1

By putting all the values in above equation, we get

M_1 (1.67)^2=83.8 (1)^2

M_1 (2.7889)=83.8

M_1 = \frac{83.8}{2.7889}=30.04g/mol

Therefore the molor mass of Hydrocarbon X is = 30.04g/mol